Lucifer February Com Night 9 February 14-15 2009

Observers: N.Ageorges, W.Seifert, V.Knierem (LUCIFER), DThompson (@LBT)
Telescope Operator: W.Wack
SW Support: CBiddick until midnight, on call thereafter.
Telescope Support: J.Rousseau

Summary:

  • HP3 continues to be a problem when M1 ventilation on (worked most of the night with it off)
  • Checked AGW transformations, still work to do.
  • M1 panic because of the lack of ventilation (c00 forces got too high).

Details - Sunset 18:11 Sunrise 7:09 LST@midnight 9:15:12

Update on HP3 Seems to only fail with mirror ventilation on?

Update on Rotator/Cable chain drive Operable. What was the fix? (see Issue #: 2017)

At sunset - mostly cloudy, thin cirrus, cleared off fairly quickly.

18:27 Starting series of closed-dome presets in TRACK mode to watch telescope collimation/HP3
First series executed with mirror ventilation off.
EL collim type Target
83 OK Preset WT10_297
83 OK Offset WT10_297
70 OK Preset WT10_246
70 OK Offset WT10_246
43 OK Preset ACT0157
43 OK Offset ACT0157
32 OK Preset WT10_123
32 OK Dither WT10_123
43 OK Preset ACT0157
43 OK Dither ACT0157
69 OK Preset WT10_246
69 OK Dither WT10_246
79 OK Preset WT10_297
79 OK Dither WT10_297

So, 7 presets and a total of 46 recollimations without complaint from HP3.

Offsets were all 200 arcsec E and N. The dither script executes 9 offsets (600-1200 arcsec) with a sleep 5 between, except for the first one that only went 200-300 arcsec dithers).

18:59 Opening dome to repeat above tests when open and with mirror ventilation on (turned on at 19:03). Tair=-7.3 Tmirror=-5.5 RH=64.0 Dewpt=-13 Wind 10 m/s with gusts to 15.

19:13 Mirror ventilation has now been on for 10 minutes. Starting repeat observations. Tests started from EL~5 and manually collimated to prevent the problem of a preset from the stow position.

EL collim type Target
31 OK Preset WT10_123
31 OK Dither WT10_123
41 OK Preset ACT0157
41 OK Dither ACT0157
65 OK Preset WT10_246
65 OK Dither WT10_246
73 OK Preset WT10_297
72 OK Dither WT10_297
63 OK Preset WT10_246
62 OK Dither WT10_246
39 OK Preset ACT0157
39 OK Dither ACT0157
28 OK Preset WT10_123
28 OK Dither WT10_123

7 presets and 70 total recollimations of M1 - 15 Failed to make tolerance on chase (21%). At:
19:16:02.958, 19:21:02.032, 19:22:27.378, 19:26:29.421, 19:26:56.125, 19:27:24.388, 19:32:22.752, 19:33:22.215, 19:37:48.797, 19:38:15.096, 19:38:43.333, 19:42:21.186, 19:43:43.295, 19:45:07.305, 19:47:35.167

Partial repeat above with mirror ventilation off.

EL collim type Target
28 OK Preset WT10_123
28 OK Dither WT10_123
37 OK Preset ACT0157
66 OK Preset WT10_297
66 OK Dither WT10_297

3 presets and 21 total recollimations of M1 - None failed to make tolerance on chase (0%).

For the 105 TRACK mode offsets that were successful in the sample I was given, I find:
mean stddev median min max max-1
13.95 2.96 13.97 8.67 25.56 20.44

20:05 Interrupting tests to go to zenith for LUCIFER to work a balky mask back into storage.

20:48 Ready to send presets, LUCIFER driving. Preset started 20:49:23.281, TRACK mode to adjust pointing.
IE changed from -31.0 to +84.68, CA from 10.0 to -33.8

21:02 GCS failed to receive a full guide cam image. First occurrence at 20:59:03. GCS restarted.

Tair=-6.2, Tmirror=-5.4 RH=17.3%, Dewpoint=-29.2

21:02:55.221 Preset in active more successful. WFS rms: 3283, 2228, 2030, 1418, 949, 323, 184

21:10:20.503 Preset samer source, off-axis guide star successful. WFS rms: 344, 272. Started script on LUCIFER. Images: luci_20090214_agw_0001 ... 0006 (useful to check offsets and orientation). EL~52. Repeating script at different position angles.

21:59 Mirror ventilation switched on.

22:07 Switched back to internal GCS transformation

22:14 Mirror ventilation back off because there was a chase failure at 22:11:04.766 (12 minutes after switching it on)

22:15:52.950 on - more offsetting tests. Seeing degrading.

Note: we have LEFTZEROPOINT = -187.651 still from Jan 17 measurements.

AGW transformations found to be "not quite there yet".

Torsten: there was a negative time for "processing overheads" in the log at Feb 15 01:39:29.

02:06:27.986 Preset failed for lack of guide star. Previous preset had star near left edge of acquisition image. Temp has gone up 5C since the start of the night (now Tair=-2.6C, Tmirror=-5.6 and not warming fast because the ventilation is still off).

03:50:07.195 Preset alert for left LUCIFER mode ACTIVE: couldn't move probe to position (SFP) (x,y,z): 4.58752, 123.4, 23.5014 mm. Like yesterday, this seems to be right in the middle of the patrol field. Why are we getting this failed message?
Sun Feb 15 03:54:12.482 2009  Preset command started for left LUCIFER mode ACTIVE from IIF
Sun Feb 15 03:54:21.908 2009  Preset alert for left LUCIFER mode ACTIVE: couldn''t move probe to position (SFP) -0.0407854, -0.41031, focus: 30.9873 :AGw command status: AGW1: Motor already in motion
Sun Feb 15 03:54:21.950 2009  Preset alert for left LUCIFER mode ACTIVE: couldn''t move probe to position (SFP) -0.0407854, -0.41031, focus: 30.9873 :AGw command status: AGW1: Motor already in motionerror: setProbePosition: couldn''t move probe to position (SFP) (x,y,z): -0.0407854, -0.41031, 30.9873 mm.
Sun Feb 15 03:54:21.987 2009  Preset alert for left LUCIFER mode ACTIVE: guidestar list successfully changed but probe couldn''t be moved to position.
Sun Feb 15 03:54:22.011 2009  Preset command failed for left LUCIFER mode ACTIVE from IIF

Nancy noticed that it said the motor was still moving. Not sure why it says this, but just in case something odd is going on we are homing the motors on the AGW unit. Subsequent preset (on the same source) successful.

Fairly quiet for some time, until:

04:20 Hit -2.6mm limit on M1 Y.. Clearing active optics and such, restarting pointing and collimation as with the start of the night because we are 5C warmer. IE changed from 79.7 to 99.86, CA changed from -27.7 to -42.22. On-axis collimation, WFS rms: 1987, 2058, 1422 hit Y limit on M1 again. Added +1mm to M1 and M2 Y global offsets and resent preset, WFS rms = Mirror panicked. Should have cleared c00 forces. The c00 plot looked like there was strong spherical on M1. Ventilation turned on and we cleared out c00 and active optics then recollimated.

Ran with mirror ventilation on for the rest of the night with only a couple of chase tolerance problems.

06:25:49 Sometime after this time (this is the last timestamp in the GCS log) PCS NaN problem. Restarted PCS and readied axes, preset did slew to position, maybe at 06:32:01, but:

06:35:49.816 Preset: Chase failed to make tolerance.

06:45 Closing dome. Tair=-3.2 Tmirror=-3.9 RH=4% Dewpoint=-39.9 Wind a steady 10m/s.

-- DavidThompson - 14 Feb 2009
Topic revision: r7 - 15 Feb 2009, DaveAshby
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback